Excercise1-7,1-8,
paad during the made by the owners Exercise 1-7 The Accounting Equation For cach of the following cases, fill in the blank with the appropriate dollar amount. LOS 3 Case 3 $40.00015.000 20000 20.000 Total assets, end of period Tocal Sabilities, end of period Capital stock, end of period Retained earnings, beginning of period Net income for the period Dividends for the period 75,000 $50000 0.000 10,000 15.000 5,000 6.00010000 9000 8.000 7,000 20,000 10,000 20,000 3.000 8.000 Exercise 1-8 Classification of Financial Statement Items LOS Classify each of the following items according to (1) whether it belongs on the income statement EXAMPLE 1-3,1-4,1-5 (IS) or balance shect (BS) and (2) whether it is a revenue (R), expense (E), asset (A), liability (I.), or stockholders equity (SE) item. Item Appears on the Classified as BS Example: Cash 1. Salaries expense 2. Equipment 3. Accounts payable 4 Membership fees earned 5. Capital stock 6. Accounts receivable 7. Buildings 8. Advertising expense 9. Retained earmings LOS EXAMPLE 1-5 Exercise 1-9 Classification of Financial Statement Items Regal Entertainment Group operates the largest chain of movie theaters in the U.S. Classify each of the following items found in the companys financial statements included in the Form 10-K for the year ended December 31, 2015, according to (1) whether it belongs on the income state- ment (IS) or balance sheet (BS) and (2) whether it is a revenue (R), expense (E), asset (A),liability (L), or stockholders equity (SE) item. REAL WOL Appears on the 8S Classified as A. Item Example: Cash and cash equivalents 1. Trade and other receivables, net 2. Clss A common stock 3. Inventories 5. Cost of concessions 6. Equipment 7. Accounts payable 8. Retained earnings 9. Interest expense, net 10. Long-term debt, less current portion
Answer #1

Answer -

EXERCISE 1-7 - THE ACCOUNTING EQUATION

Case 1 : Total liabilities, end of period = $ 9,000.

Working Notes :

1).Retained earnings, end of period =

Beginning balance + Net Income - Dividend

=15,000 + 8,000 - 2,000 = $ 21,000

2).Total liabilities, end of period =

Total Assets, end of period - (Capital stock, end of period + Retained earnings, end of period)

=$ 40,000 - (10,000 + 21,000) = $ 9,000

Case 2 : Total Asset, end of period = $ 34,000.

Working Notes -

1). Retained earnings, end of period =

Beginning balance + Net Income - Dividend paid

= 8,000 + 7,000 - 1,000 = $ 14,000

2). Total Assets, end of period =

Total liabilities, end of period + Capital stock, end of period + Retained earnings, end of period

= 15,000 + 5,000 + 14,000 = $ 34,000

Case 3 : Net income for the period = $ 23,000.

Working Notes

1). Retained earnings, end of period =

Total Assets, end of period  - ( Total liabilities, end of period + Capital stock, end of period )

= $ 75,000 - ( 25,000 + 20,000) = $ 30,000

2). Net income for the period =

Retained Earnings, end of period - Retained earnings, beginning of period + Dividend paid

= $ 30,000 - 10,000 + 3,000 = $ 23,000

Case 4 : Dividend for the period = $ 4,000.

Working Notes -

1). Retained earnings, end of period =

Total Assets - ( Total liabilities + Capital stock )

= $50,000 - ( 10,000 + 15,000) = $ 25,000

2). Dividend for the year =

Retained earnings, beginning of period - Net income + Retained earnings, end of period

= $ 20,000 + 9,000 - 25,000 = $ 4,000

EXERCISE 1-8 - CLASSIFICATION OF FINANCIAL STATEMENT ITEMS

Item Appears on the Classified as
Example: Cash BS A
1. Salaries Expense IS E
2. Equipment BS A
3. Accounts payable BS L
4. Membership fees earned IS R
5. Capital stock BS SE
6. Accounts receivable BS A
7. Building BS A
8. Advertising Expense IS E
9. Retained earnings BS SE

EXERCISE 1-9 CLASSIFICATION OF FINANCIAL STATEMENT ITEMS

Item Appears on the Classified as
1. Trade and other receivables, net BS A
2. Class A Common stock BS SE
3. Inventories BS A
4. Admissions IS R
5. Cost of concessions IS E
6. Equipment BS A
7. Accounts payable BS L
8. Retained earnings BS SE
9. Interest expense, net IS E
10. Long term debt, less current portion BS L